Engineering/D-Night 20180522 (UT)

Night Plan:

  • Sunset 19:21
  • Moon 51% - sets 01:22
  • Twilight 20:52, 03:40
  • Night Center 00:16
  • Sunrise 05:11

  • D-time observations at start
    • DONE Q1330+011, LUCI, zJ+HK, 1 hour, start of night (Olga).
    • DONE Local Mergers - LUCI-ESM WAS49 K-band binocular (Barry)
    • Local Mergers - LUCI (Barry) - punted science for a later date
    • DONE Request from AO Group for more ESM on-sky-science type observations - script prepared for NGC 3921 (Ks binocular, and Fraternal LUCI-1 J LUCI-2 H - total time 30min w/overheads for all three filters) - also tried M51
  • Arcetri AO Testing - 4 hours requested - start 11:40, end 03:40
    • Test loop stability
    • TT LUT?
    • Seeing < 1.2" preferred, can accomodate seeing up to 1.5"
  • In case of bad seeing (> 1.5"):
    • DONE - 1 HK target (D-time) Local Mergers - Near-IR spectroscopy (zJHK) seeing up to ~2", several mags of extinction - LUCI-Bino (Barry)
    • (D-time) H-band spectral library, seeing >>2" - many mags extinction - LUCI-Bino G210-H (Barry)
    • (E-time) TMS Active with LBC (John H. and Andrew) (seeing <2", up to 3 mags clouds, target flexible)
    • (E-time) LUCI+MODS Binocular preset tests ( IT 6343) - has not been addressed in over a year, but the IT suggests the fix needs to be tested (John H. and Barry)
    • (E-time) LBC+MODS pseudo-monocular test ( IT 5715) - verify that LBC will dither without streaking in exposures (John H. and Barry)
    • (E-time) LBC - WRS testing on tech chips


OSA: G. Bechetti

ISp: J. Morris

ISA: B. Rothberg

AO: D. Miller

Mtn Manager: M. Wagner

UT 01:58 - Took 2 test darks on each side L1 and L2 (0001-0002)

UT 02:03 - D. Miller and G. Bechetti running through AO checkout

UT 02:15 - Opening enclosure

UT 02:31 - Moving to pointing and colimation star checks (too bright still)

UT 02:43 - sending active preset to collimate

UT 02:48 - plan will be to take some quick images with both LUCIs using N30 camera and mask in turnout before heading to first science target

UT 02:52 - first WFS AGW images shows ~0.8" seeing - looks promising for the evening

UT 02:58 - 1.05" on DIMM

UT 03:11 - taking AO checkout images with LUCI via script

===========================================

Q1330+011 - Olga

UT 03:12 - Moving to telluric for science data

UT 03:38 - Moving to science target

UT 03:44 - acquiring QSO in slits

UT 03:57 - taking science spectra

UT 04:14 - DIMM 0.75" AGWS 0.67" 0.75"

UT 04:44 - Finished QSO

===========================================

WAS49 - ESM test - Barry

UT 04:45 - moving to Standard Star P266-C (Ks) L1#0040-0043 ; L2# 0042-0045

UT 04:48 - 0.99" DIMM 0.7-0.9" on AGWs - Airmass 1.039

UT 04:54 - Send preset for ESM Binocular WAS 49

UT 05:02 - starting exposures for ESM - 0.65"-0.68" on guiders L1#0044-0068 ; L2#0046-0070 (first couple or three images might

UT 05:04 - very long wait for offset on SX side

UT 05:10 - corrections a bit better on SX ~0.36"-0.40, after L2#0050 will adjust focus a bit on DX to imrpove image quality.

UT 05:11 - getting windy, rocking the DX side (not in a good way!)

UT 05:12 - from 55.2 to 55.5mm adjustment on DX to attempt to improve focus

UT 05:15 - improved so L1 and L2 ~ similar FWHMs (thanks Doug and Greg!) (comparing L1#0049 vs L2#0051)

UT 05:20 - DX side degrading a little bit again (wind?)

UT 05:23 - winds hitting 18 m/s

UT 05:26 - DIMM seeing ~1.2" - airmass 1.09

UT 05:39 - finished 20exposures, going back to start of script to get a few more

UT 05:44 - DX waiting for IIF....success after 1 minute waiting

UT 05:49 - finished WAS49 - airmass 1.140

UT 05:50 - moving back to P266-C - airmass 1.137

UT 05:59 - Finished P266-C, moving to NGC 3921 as another ESM test (since AGW-1 is coming off this is our last chance until 2018B to do any ESM characterization)

=================================================

ESM testing

UT 06:01 - DIMM 1.26" - airmass 1.25

UT 06:02 - keeping an eye on cloud in the NE

UT 06:06 - still waiting for collimation

UT 06:07 - starting first exposure NGC 3921 Ks ESM L1#0073 -0081 ; L2#0075-0083

UT 06:13 - 51 elevation, getting 7-8 pixel FWHM on both sides, but less wind

UT 06:19 - increased gains on LUCI-1 (starting with #0080)

UT 06:20 - keeping preset, manually configuring L1 (J) L2 (H) - skipping to 4th item in script then will back track L1# 0082-0088 ; L2# 0084-0090

Dithers:

-15,15 on

15,-15 on

-315,-285 off

-300,-285 off

-15,-15 on

15, 15 on

-315,-300 off

UT 06:37 - Moving to M51 - Ks AORef star 15 - accidentally clicked ARGOS -GS which is track - remade script with LUCI-GS

UT 06:41 - resent preset - failed due to interrupt call

UT 06:42 - sending preset again

UT 06:43 - L1#0089-0097 ; L2#0091-0099

UT 06:49 - starting first exposure

UT 06:58 - 1.6" in direction of target, 1.1" on images

UT 07:00 - running with Br Gamma on LUCI-1 and H on LUCI-2 L1#0098-0106; L2#0100-0108 ; 1.65" on AO GUIs

UT 07:21 - completed M51 ESM test

=================================================

UT 07:21 - handing over to Arcetri....with 1.46" on the DIMM (good luck!)

UT 07:23 - putting N30/N30fieldstop in but going to zenith and parking rotator first L1-341 L2-210

UT 07:25 - LUCI-1 Camera went to Unknown - initializing on IMGUI, that didnt' work had to use Engineering panel to initialize and move

UT 07:31 - MOS successful, OH1060/z on LUCI-1, FeII on LUCI-2

================================================

UT 09:07 - returning to LBTO for poor conditon program - Barry - Local Mergers nearIR

UT 09:08 - Slewing to Telluric for NGC 6598 (HIP 82884)

UT 09:11 - Guider 1.1"

UT 09:19 - offset +0.2" DETX on the alignment slit

UT 09:20 - HIP 82884 G200 HK - L1#0112-0114 ; L2#0123-0126 Airmass 1.093

UT 09:21 - error tilt grating LUCI-1 - so had to manually reset and skip to 2nd line in spectroscopy (L1 will be missing one exposure)

UT 09:27 - HIP 82884 G200 zJ (loading zJ and just continuing, alignment should be fine) L1#0115-0118 ; L2#0127-0130

UT 09:29 - DIMM 1.48" guiders 1.52"

UT 09:35 - moving to science field NGC 6598 - airmass 1.24

UT 09:42 - LUCI-1 RTD is super slow for alignment - subtracted image taking forever

UT 09:52 - stretch is really annoying on RTD - think need to nudge the L2 over 2 pixels - seeing improving ~1.1-1.2"

UT 09:56 - starting science L1#0123-0132 ; L2#0136-0145

UT 10:19 - DIMM 1.2-1.3" but AGWs show ~2"

UT 10:26 - DIMM hit ~1.2-1.3" - showed one 0.95" point - got Doug M. to check if conditions are usable for any AO

UT 10:30 - Local Merger near-IR program done for now

==========================

UT 10:31 - Configuring LUCI-1 for N30/N30_fieldstop, put blind, mirror no flexure LUCI-2

UT 10:36 - Setup complete for LUCI-1, moving to AO star then will check mask alignment

UT 10:37 - DIMM now showing 1-1.1"

UT 10:40 - checking N30 fieldstop alignment on LUCI-1

UT 10:42 - alignment done, needed 2 moves, switching to clear/FeII L1

UT 10:49 - sub-windowed 128x128 LUCI-1 - arcetri starting collection

UT 11:46 - Arcetri still taking data on SX, DX idle

UT 11:52 - loops opened....AdSec arbitrator being restarted (catch those last photons before LUCI-1 smells like BBQ...)

UT 11:55 - done with LUCI-1, putting in safe state

UT 12:01 - primary mirrors are no longer moving - the telescope is dropping hints....

UT 12:04 - telling Arcetri we are closing up (sometimes subtley is a fine art...)

UT 12:05 - Closing Enclosure - will take LUCI cals afterwards

=================================================

UT 12:15 - starting LUCI-Bino cals

==================

0.5" HK G200 LUCI1-LUCI2 Arcs/Flats for Local Mergers near-IR program (Barry)

L1#0152-0156 HK flats lights off, L1#0157-0161 HK flats lights on, L1#0162-0163 Arcs lights off, 0164-0165 Argon, 0166-0167 Xenon, 0168-0169 Neon

L2#0146-0150 HK flats lights off, L2#0151-0155 HK flats lights on, L2#0156-0157 Arcs lights off, 0158-0159 Argon, 0160-0161 Xenon, 0162-0163 Neon

================================

1" HK G200 LUCI1-LUCI2 Arcs/Flats for QSO program (Olga)

L1# 0170-0174 HK flats lights off, 0175-0179 HK flats lights on, 0180-0181 Neon, 0182-0183 arcs lights off, 0184-0185 Argon, 0186-0187 arcs lights off, 0188-0189 Xenon, 0190-0191 arcs lights off

L2# 0164-0168 HK flats lights off, 0169-0173 HK flats lights on, 0174-0175 Neon, 0176-0177 arcs lights off, 0178-0179 Argon, 0180-0181 arcs lights off, 0182-0183 Xenon, 0184-0185 arcs lights off

UT 12:48 - 1" zJ G200 LUCI1-LUCI2 Arcs/Flats for QSO program (Olga)

L1#0192-0196 zJ flats lights off, 0197-0201 zJ flats lights on, 0202-0203 Neon, 0204-0205 arcs lights off, 0206-0207 Argon, 0208-0209 arcs lights off

L2#0186-0190 zJ flats lights off, 0191-0195 zJ flats lights on, 0196-0197 Neon, 0198-0199 arcs lights off, 0200-0201 Argon, 0202-0203 arcs lights off

=================================

Imaging flats for ESM (Ks for L1/L2, J,Br gamma for L1, H for L2) (Barry)

UT 13:07 - configuring for Imaging flats

L1#0210-0214 Ks flats lights off, 0215-0219 Ks flats lights on, 0220-0224 J flats lights off, 0225-0229 J flats lights on, 0230-0234 Br Gamma lights off, 0235-0239 Br Gamma lights on

L2#0204-0208 Ks flats lights off, 0209-0213 Ks flats ligths on, 0214-0218 J flats lights off, 0219-0223 J flats lights on

UT 13:18 - LUCI-2 put into safe mode

UT 13:23 - LUCI-1 put into safe mode


 
Topic revision: r10 - 22 May 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback